how do you find the slope parallel ro the graph
of 3x+4y=12

Answer by user_dude2008(1862)

3x+4y=12
4y=12-3x
4y=-3x+12
y=(-3x)/4+(12)/4
y=-3/4x+3
Slope is -3/4